OpenStudy (anonymous):

Please help Lupita rides a taxi that charges a flat rate of $6.75 plus $3.20 per mile. If the taxi charges Lupita $40.03 in total for her trip, how many miles is her ride? Enter your answer in the box as a decimal to the nearest tenth of a mile. Is it 12.5?

OpenStudy (anonymous):

@Lovelarap

OpenStudy (anonymous):

Aw, baby Skipper lol

OpenStudy (anonymous):

I think... oops

OpenStudy (jiteshmeghwal9):

i am rico

OpenStudy (anonymous):

Oh yeah, still cute haha

OpenStudy (anonymous):

Wait did u draw that yourself Lovelrap? in your profile pic

OpenStudy (anonymous):

Ya i did. :)

OpenStudy (anonymous):

Oh cool, I like to draw too :DD

OpenStudy (anonymous):

I have another .

OpenStudy (anonymous):

Answer or drawing

OpenStudy (anonymous):

Kinda both

OpenStudy (anonymous):

Oh, well how did you get your answer?

OpenStudy (anonymous):

I got 18.98 (rounded)

OpenStudy (anonymous):

oh, did you add the 6.70?

OpenStudy (anonymous):

6.75 I mean

OpenStudy (anonymous):

lovelrap?

OpenStudy (anonymous):

No, i multiplyed 3.20 by 40.30 then divided by6.75

OpenStudy (anonymous):

Sorry my tablet is slow and kinda laging :(

OpenStudy (anonymous):

oh, wouldn't you multiply though?

OpenStudy (anonymous):

i mean divide

OpenStudy (anonymous):

Well ya are there any choices?

OpenStudy (anonymous):

I just divided 3.20 and 40.50, is that wrong?

OpenStudy (ella4184):

X=number of miles f(x)= money spent in $ f(x)=3.2x+6.75 if; f(x)=40.03 then: 40.03=3.2x+6.75 3.2x=40.03-6.75 3.2x=33.28 x=33.28/3.2=10.4 Solution: 10,4 miles

OpenStudy (ella4184):

Hope this helps! :)

OpenStudy (anonymous):

why did you subtract and switch 40.03 and 3.2x?

OpenStudy (anonymous):

wait nvm I get it

OpenStudy (ella4184):

Good :)

OpenStudy (anonymous):

Can someone help in history?

OpenStudy (anonymous):

Thanks both @Lovelarap and @Ella4184

OpenStudy (ella4184):

I can try. What is your question?

OpenStudy (ella4184):

@Lovelarap